# Roadmap for electronic structure, anharmonicity, and electron-phonon calculations in locally disordered inorganic and hybrid halide perovskites

### v1

@@ I Introduction
Halide perovskites, either layered, bulk, nanocrystals, or heterostructures are attracting immense interest due to their impressive power conversion efficiencies in optoelectronics, with new candidates emerging constantly Kojima et al. (2009); Snaith (2013); Filip et al. (2014); Protesescu et al. (2015); Volonakis et al. (2016); Eperon et al. (2016); Tsai et al. (2016); Brenner et al. (2016); Volonakis et al. (2017); Mao et al. (2018); Hoffman et al. (2019); Liu et al. (2020); Marchenko et al. (2020); Dey et al. (2021); Ahmadi et al. (2021); Jiang et al. (2022); Jacobsson et al. (2021); Unger and Jacobsson (2022); Sidhik et al. (2022); Metcalf et al. (2023); Sidhik et al. (2024); Filip and Leppert (2024). Several key challenges in the field of halide perovskites remain partially unexplored and continue to be active areas of research. These include gaining a deeper understanding of (i) the complex anharmonic potential energy surface responsible for the formation of local structure/disorder Even et al. (2016); Beecher et al. (2016); Marronnier et al. (2018); Zhao et al. (2020); Fabini et al. (2020); Hehlen et al. (2022); Dirin et al. (2023); Zacharias et al. (2023a); Weadock et al. (2023); Diez-Cabanes et al. (2023); Balvanz et al. (2024); Caicedo-Dávila et al. (2024); Dubajic et al. (2025), anharmonic lattice vibrations Quarti et al. (2015); Carignano et al. (2017); Yaffe et al. (2017); Katan et al. (2018); Marronnier et al. (2018); Fu et al. (2018); Ferreira et al. (2020); Zacharias et al. (2023a, b), and phase stability Marronnier et al. (2017); Sutton et al. (2018); Marronnier et al. (2018); Doherty et al. (2021); He et al. (2024), and (ii) anharmonic carrier-phonon coupling and its role to the electronic structure renormalization Patrick et al. (2015); Quarti et al. (2016); Ning et al. (2022); Zacharias et al. (2023a); Seidl et al. (2023), temperature-dependent optical absorption, photoluminescence (PL) Maculan et al. (2015); Milot et al. (2015); Wright et al. (2016); Kontos et al. (2018); Hsu et al. (2019); Mannino et al. (2020); Kahmann et al. (2020); Chen et al. (2020); Peters et al. (2022), carrier mobilities Stoumpos et al. (2013); Milot et al. (2015); Parrott et al. (2016); Poncé et al. (2019); Irvine et al. (2021); Cucco et al. (2024), polaron formation Zhu and Podzorov (2015); Frost (2017); Miyata et al. (2017); Bretschneider et al. (2018); Cinquanta et al. (2019); Ghosh et al. (2020); Österbacka et al. (2020); Srimath Kandada and Silva (2020); Puppin et al. (2020); Cannelli et al. (2021); Guzelturk et al. (2021); Buizza and Herz (2021); Schilcher et al. (2021); Zhang et al. (2021); Yue et al. (2023); Lafuente-Bartolome et al. (2024); Baranowski et al. (2024), and ultrafast non-equillibrium dynamics Cinquanta et al. (2019); Zhang et al. (2023); Seiler et al. (2023); Biswas et al. (2024).
@@
Ab initio electronic structure and electron-phonon calculations for halide perovskites typically utilize a simple monomorphous network, which can be described by a minimal unit cell Huang and Lambrecht (2013); Even et al. (2013a, b); Amat et al. (2014); Umari et al. (2014); Mosconi et al. (2016); Bokdam et al. (2016); Saidi et al. (2016); Whalley et al. (2017); Mao et al. (2018); Traoré et al. (2019); Poncé et al. (2019); Ghaithan et al. (2020); Wang et al. (2020b); Ning et al. (2022); Muhammad et al. (2022). However, this structure is dynamically unstable, corresponding to a local maximum on the PES [Fig. 1]. As a result, density-functional theory (DFT) calculations in halide perovskites based on monomorphous cells fail to capture the physics related to the more stable polymorphous networks, which are known to significantly influence the electronic structure Zhao et al. (2020). In Ref. [Zacharias et al., 2023a], relying on the anharmonic special displacement method (ASDM) Zacharias and Giustino (2016, 2020); Zacharias et al. (2023b), we have introduced a bottom-up ab initio approach for the exploration of polymorphous structures and a unified treatment of the complex interplay between local disorder, anharmonicity, and electron-phonon coupling. We have demonstrated using the case of inorganic Cs-based halide perovskites that local disorder plays a key role to accurately describe strongly overdamped vibrational dynamics, anharmonic electron-phonon coupling, and temperature-dependent band gaps. In Ref. [Zacharias et al., 2025], we have generalized and applied our approach in a high-throughput framework to the computationally more demanding and technologically more significant case of cubic hybrid halide perovskites.
@@
In this paper, we provide a detailed description of our approach and conduct a comprehensive comparative analysis of 12 materials (4 FA-, 4 MA-, and 4 Cs-based compounds) through large-scale calculations for 120 polymorphous structures (i.e. 10 structures each) using supercells. We obtain the following results: (i) We introduce the concept of reference structures for hybrid halide perovskites, different than the monomorphous structures, where disorder is accounted initially only for the molecules. The reference structures enable the separation of the effects arising from the distinct dynamics of organic molecules and facilitate systematic comparisons with the electronic structure of the corresponding fully polymorphous states. (ii) We quantify local disorder in terms of B-X bond length and B-X-B bond angle distortions and demonstrate strong correlations with the increase in band gap and effective masses induced by positional polymorphism. (iii) We demonstrate that the extent of local disorder is influenced by the size of the A-site cation, with FA-based compounds exhibiting the least positional polymorphism, followed by MA-based compounds, and Cs-based compounds displaying the highest degree of polymorphism. (iv) We compute anharmonic phonons at finite temperatures within the self-consistent phonon theory based on reference structures, and present phonon quasiparticle dispersions for all compounds. Our results compare well with diffuse scattering maps and phonon frequencies measured by various experimental techniques. We also discuss the role of A-site cations and X-site anions to transverse optical (TO) and longitudinal optical (LO) phonon splitting. (v) We calculate the free energies and Boltzmann-weighted averages of temperature-dependent band gaps of all cubic compounds using both reference and polymorphous structures, highlighting the critical role of local disorder in achieving accurate anharmonic electron-phonon coupling calculations, which result in unprecedented agreement with experiments Milot et al. (2015); Kontos et al. (2018); Parrott et al. (2016); Sutton et al. (2018); Hsu et al. (2019); Mannino et al. (2020); Chen et al. (2020); Kahmann et al. (2020); Peters et al. (2022); López et al. (2025). (vi) We distinguish the contributions of thermal expansion and electron-phonon coupling to band gap renormalization, demonstrating that they do not contribute equally, contrary to common assumptions Francisco-López et al. (2019); Rubino et al. (2020). Overall, electron-phonon coupling dominates for Cs-based and MA-based compounds, reaching as high as 97% while for FA-based compounds thermal expansion plays more important role. For locally disordered Sn-based compounds, we find an enhanced thermal expansion contribution to the band gap renormalization due to the enhanced lone pair activity, promoting Sn-I-Sn bond angle bending under lattice expansion. (vii) We calculate temperature-dependent hole and electron effective masses of all cubic compounds using both reference and polymorphous structures. The effective mass enhancement due to polymorphism and electron-phonon coupling reach as high as on order of magnitude, improving agreement between calculated and measured reduced effective masses. (viii) We show that, starting from maximally symmetrized reference structures, even one polymorphous configuration can be enough to calculate anharmonic electron-phonon properties in some systems. (ix) We show that positional polymorphism leads to a continuum of strongly overdamped vibrations and the breakdown of the phonon quasiparticle picture obtained by standard approaches to anharmonicity applied to monomorphous or reference cells Patrick et al. (2015); Ning et al. (2022); Zacharias et al. (2023b).
@@
We organize the manuscript as follows: In Sec. II we introduce the theory underpinning our electron-phonon coupling, temperature-dependent electron energies, polymorphous structure, and anharmonic phonon calculations via the ASDM. Section III described our main methodology for computing anharmonic electron-phonon properties in locally disordered halide perovskites, starting from reference structures. In the same section, we show the total energy lowering obtained for locally disordered structures with respect to reference structures and present strong correlations of the band gap widening and effective masses enhancements with bond length and angle variations. We also demonstrate our methodology for computing anharmonic phonon quasiparticle dispersions, and free energies, as well as Boltzmann-weighted averages of temperature-dependent band gaps and effective masses of MAPbBr3 and FAPbI3. In Sec. IV, we provide a case-by-case analysis of all 12 halide perovskites, showing band gaps at the DFT and hybrid functional levels, DFT effective masses, analyzing anharmonic phonon dispersions and TO,LO phonon energies, and temperature-dependent band gaps and effective masses calculated within the ASDM. Where available we compare our results with state-of-the-art $GW$ calculations. In the same section, we analyze the electron-phonon and thermal expansion contributions to the band gap renormalization for each material. In Sec. V, we present a general comparison and discussion of our results focusing on electronic structure modifications due to local disorder, LO-TO splitting, the breakdown of the phonon quasiparticle picture, and thermal induced renormalization. Section VII summarizes our key results and conclusions as well as provides avenues of future work.
@@ II.1 Electron-phonon coupling
$g_{nm\nu}=\sqrt{\frac{\hbar}{2M_{0}\omega_{\nu}}}\braket{\psi_{n}}{\,\partial U^{\{\tau\}}_{\rm KS}/\partial x_{\nu}\,}{\psi_{m}}$ (1)
@@
Figure 2: (a) Schematic representation of the transitions of the PES from the monomorphous to the polymorphous, and from the polymorphous to the reference structure, indicated by ball and stick models of cubic FAPbI3. (b) DFT total energy lowering per formula unit ($\Delta U$) of polymorphous cubic halide perovskites relative to their reference structures, evaluated using Eq. (11), assuming $T=300$ K, and 10 configurations in $2\times 2\times 2$ supercells for each system. Horizontal gray, red, and light-blue lines indicate the mean $\Delta U$ of FA-, MA-, and Cs-based compounds, respectively. (c-h) Scatter plots showing the dependence of the band gap and effective mass increase due to polymorphism on the average B-X bond length and B-X-B bond angle variations. The 120 data points correspond to 10 polymorphous configurations generated for 4 FA (black), 4 MA (red), and 4 Cs (blue) -based compounds. The straight thick lines represent the least squares regression fit to the data, with the slopes indicated on each plot. The shaded regions represent three standard deviations on either side of the lines. All calculations refer to the DFT-PBEsol approximation. Our calculations for the band gaps and effective masses include spin-orbit coupling (SOC) effects.
@@ II.2 Temperature-dependent electron energies
where the first term inside the square brackets represents the Fan-Migdal contribution and $h_{n\nu\nu}=\hbar/2M_{0}\omega_{\nu}\braket{\psi_{n}}{\,\partial^{2}U^{\{\tau\}}_{\rm KS}/{\partial x_{\nu}^{2}}\,}{\psi_{n}}$ represents the Debye-Waller contribution to electron-phonon renormalized band structures Giustino (2017). We note that the electron-phonon matrix elements [Eq. (1)] entering Eq. (4), when computed within the harmonic approximation, may be ill-defined if the system exhibits imaginary phonon modes. This is usually an artifact of applying the harmonic approximation to an inherently anharmonic system. This underscores the need to explicitly consider the system’s anharmonicity, as discussed in Secs. II.3 and II.4. Another source of anharmonicity affecting the electronic structure at finite temperatures is thermal lattice expansion, which is treated separately in this work, as discussed in Ref. [Zacharias et al., 2025] and Sec. VI.6.
@@ II.3 Polymorphous structure exploration
Using the example of CsPbI3 for a $2\times 2\times 2$ supercell, we tested the values $x_{\nu}=\{0.05,0.5,1,2,3,4,5\}$  to initially displace the atoms of the reference structure using Eq. (5), followed by structural relaxation. For $x_{\nu}=\{0.5,1,2\}$, the system reached a polymorphous structure, resulting in a total energy lowering of 121 meV relative to the reference structure. For $x_{\nu}=\{3,4,5\}$, the system converged to a saddle point with a total energy reduction of 108 meV, whereas for $x_{\nu}=0.05$, the system returned back to the reference structure.
@@
Here, the summation is performed over stable phonons modes with $\omega^{2}_{\nu}>0$ and $S_{\nu}$ represents signs associated with each eigenvector. These signs are assigned combinatorially by ZG.x of EPW Lee et al. (2023) to ensure that the anisotropic displacement parameters are closely reproduced Zacharias and Giustino (2020). Using CsPbI3, we tested ZG displacements for $T=0$ K. While the system after the escape step moves toward a less stable configuration than the reference structure, the final relaxation step leads to a polymorphous structure with the same energy lowering of 121 meV, as achieved using Eq. (5) for $x_{\nu}=\{0.5,1,2\}$. We note that for generating polymorphous structures with distinct atomic coordinates, one can employ different initial sets of ZG displacements by using various permutations of the set of signs $S_{\nu}$ or different temperatures $T$.

### v2

@@ I Introduction
Halide perovskites, either layered, bulk, nanocrystals, or heterostructures are attracting immense interest due to their impressive power conversion efficiencies in optoelectronics, with new candidates emerging constantly Kojima et al. (2009); Snaith (2013); Filip et al. (2014); Protesescu et al. (2015); Volonakis et al. (2016); Eperon et al. (2016); Tsai et al. (2016); Brenner et al. (2016); Volonakis et al. (2017); Mao et al. (2018); Hoffman et al. (2019); Liu et al. (2020); Marchenko et al. (2020); Dey et al. (2021); Ahmadi et al. (2021); Jiang et al. (2022); Jacobsson et al. (2021); Unger and Jacobsson (2022); Sidhik et al. (2022); Metcalf et al. (2023); Sidhik et al. (2024); Filip and Leppert (2024). Several key challenges in the field of halide perovskites remain partially unexplored and continue to be active areas of research. These include gaining a deeper understanding of (i) the complex anharmonic potential energy surface responsible for the formation of local structure/disorder Even et al. (2016); Beecher et al. (2016); Marronnier et al. (2018); Zhao et al. (2020); Fabini et al. (2020); Hehlen et al. (2022); Dirin et al. (2023); Zacharias et al. (2023a); Weadock et al. (2023); Diez-Cabanes et al. (2023); Balvanz et al. (2024); Caicedo-Davila et al. (2024); Dubajic et al. (2025), anharmonic lattice vibrations Quarti et al. (2015); Carignano et al. (2017); Yaffe et al. (2017); Katan et al. (2018); Marronnier et al. (2018); Fu et al. (2018); Ferreira et al. (2020); Zacharias et al. (2023a, b), and phase stability Marronnier et al. (2017); Sutton et al. (2018); Marronnier et al. (2018); Doherty et al. (2021); He et al. (2024), and (ii) anharmonic carrier-phonon coupling and its role to the electronic structure renormalization Patrick et al. (2015); Quarti et al. (2016); Ning et al. (2022); Zacharias et al. (2023a); Seidl et al. (2023), temperature-dependent optical absorption, photoluminescence (PL) Maculan et al. (2015); Milot et al. (2015); Wright et al. (2016); Kontos et al. (2018); Hsu et al. (2019); Mannino et al. (2020); Kahmann et al. (2020); Chen et al. (2020); Peters et al. (2022), carrier mobilities Stoumpos et al. (2013); Milot et al. (2015); Parrott et al. (2016); Poncé et al. (2019); Irvine et al. (2021); Cucco et al. (2024), polaron formation Zhu and Podzorov (2015); Frost (2017); Miyata et al. (2017); Bretschneider et al. (2018); Cinquanta et al. (2019); Ghosh et al. (2020); Österbacka et al. (2020); Srimath Kandada and Silva (2020); Puppin et al. (2020); Cannelli et al. (2021); Guzelturk et al. (2021); Buizza and Herz (2021); Schilcher et al. (2021); Zhang et al. (2021); Yue et al. (2023); Lafuente-Bartolome et al. (2024); Baranowski et al. (2024), and ultrafast non-equillibrium dynamics Cinquanta et al. (2019); Zhang et al. (2023); Seiler et al. (2023); Biswas et al. (2024).
@@
Ab initio electronic structure and electron-phonon calculations for halide perovskites typically utilize a simple monomorphous network, which can be described by a minimal unit cell Huang and Lambrecht (2013); Even et al. (2013a, b); Amat et al. (2014); Umari et al. (2014); Mosconi et al. (2016); Bokdam et al. (2016); Saidi et al. (2016); Whalley et al. (2017); Mao et al. (2018); Traoré et al. (2019); Poncé et al. (2019); Ghaithan et al. (2020); Wang et al. (2020b); Ning et al. (2022); Muhammad et al. (2022). However, this structure is dynamically unstable, corresponding to a local maximum on the PES [Fig. 1]. As a result, density-functional theory (DFT) calculations in halide perovskites based on monomorphous cells fail to capture the physics related to the more stable polymorphous networks, which are known to significantly influence the electronic structure Zhao et al. (2020). In Ref. [Zacharias et al., 2023a], relying on the anharmonic special displacement method (ASDM) Zacharias and Giustino (2016, 2020); Zacharias et al. (2023b), we have introduced a bottom-up ab initio approach for the exploration of polymorphous structures and a unified treatment of the complex interplay between local disorder, anharmonicity, and electron-phonon coupling. We have demonstrated using the case of inorganic Cs-based halide perovskites that local disorder plays a key role to accurately describe strongly overdamped vibrational dynamics, anharmonic electron-phonon coupling, and temperature-dependent band gaps. In Ref. [Zacharias et al., 2026], we have generalized and applied our approach in a high-throughput framework to the computationally more demanding and technologically more significant case of cubic hybrid halide perovskites.
@@
In this paper, we provide a detailed description of our approach and conduct a comprehensive comparative analysis of 12 materials (4 FA-, 4 MA-, and 4 Cs-based compounds) through large-scale calculations for 120 polymorphous structures (i.e. 10 structures each) using supercells. We obtain the following results: (i) We introduce the concept of reference structures for hybrid halide perovskites, different than the monomorphous structures, where disorder is accounted initially only for the molecules. The reference structures enable the separation of the effects arising from the distinct dynamics of organic molecules and facilitate systematic comparisons with the electronic structure of the corresponding fully polymorphous states. (ii) We quantify local disorder in terms of B-X bond length and B-X-B bond angle distortions and demonstrate strong correlations with the increase in band gap and effective masses induced by positional polymorphism. (iii) We demonstrate that the extent of local disorder is influenced by the size of the A-site cation, with FA-based compounds exhibiting the least positional polymorphism, followed by MA-based compounds, and Cs-based compounds displaying the highest degree of polymorphism. (iv) We compute anharmonic phonons at finite temperatures within the self-consistent phonon theory based on reference structures, and present phonon quasiparticle dispersions for all compounds. Our results compare well with diffuse scattering maps and phonon frequencies measured by various experimental techniques. We also discuss the role of A-site cations and X-site anions to transverse optical (TO) and longitudinal optical (LO) phonon splitting. (v) We calculate the free energies and Boltzmann-weighted averages of temperature-dependent band gaps of all cubic compounds using several reference and polymorphous structures, to account for configurational entropy effects. Our results highlight the critical role of local disorder in achieving accurate anharmonic electron-phonon coupling calculations, which result in unprecedented agreement with experiments Milot et al. (2015); Kontos et al. (2018); Parrott et al. (2016); Sutton et al. (2018); Hsu et al. (2019); Mannino et al. (2020); Chen et al. (2020); Kahmann et al. (2020); Peters et al. (2022); López et al. (2025). (vi) We distinguish the contributions of thermal expansion and electron-phonon coupling to band gap renormalization, demonstrating that they do not contribute equally, contrary to common assumptions Francisco-López et al. (2019); Rubino et al. (2020). Overall, electron-phonon coupling dominates for Cs-based and MA-based compounds, reaching as high as 97% while for FA-based compounds thermal expansion plays more important role. For locally disordered Sn-based compounds, we find an enhanced thermal expansion contribution to the band gap renormalization due to the enhanced lone pair activity, promoting Sn-I-Sn bond angle bending under lattice expansion. (vii) We calculate temperature-dependent hole and electron effective masses of all cubic compounds using both reference and polymorphous structures. The effective mass enhancement due to polymorphism and electron-phonon coupling reach as high as one order of magnitude, improving agreement between calculated and measured reduced effective masses. (viii) We show that, starting from maximally symmetrized reference structures, even one polymorphous configuration can be enough to calculate anharmonic electron-phonon properties in some systems. (ix) We show that positional polymorphism leads to a continuum of strongly overdamped vibrations and the breakdown of the phonon quasiparticle picture obtained by standard approaches to anharmonicity applied to monomorphous or reference cells Patrick et al. (2015); Ning et al. (2022); Zacharias et al. (2023b). (x) We demonstrate the role of local disorder to accurately describe the band gap variations across different phases.
@@
We organize the manuscript as follows: In Sec. II we introduce the theory underpinning our electron-phonon coupling, temperature-dependent electron energies, polymorphous structure, and anharmonic phonon calculations via the ASDM. Section III described our main methodology for computing anharmonic electron-phonon properties in locally disordered halide perovskites, starting from reference structures. In the same section, we show the total energy lowering obtained for locally disordered structures with respect to reference structures and present strong correlations of the band gap widening and effective masses enhancements with bond length and angle variations. We also demonstrate our methodology for computing anharmonic phonon quasiparticle dispersions, and free energies, as well as Boltzmann-weighted averages of temperature-dependent band gaps and effective masses of MAPbBr3 and FAPbI3. In Sec. IV, we provide a case-by-case analysis of all 12 halide perovskites, showing band gaps at the DFT and hybrid functional levels, DFT effective masses, analyzing anharmonic phonon dispersions and TO,LO phonon energies, and temperature-dependent band gaps and effective masses calculated within the ASDM. Where available we compare our results with state-of-the-art $GW$ calculations. In the same section, we analyze the electron-phonon and thermal expansion contributions to the band gap renormalization for each material. In Sec. V, we present a general comparison and discussion of our results focusing on electronic structure modifications due to local disorder, the band gap across different polymorphs, LO-TO splitting, the breakdown of the phonon quasiparticle picture, and thermal induced renormalization. Section VII summarizes our key results and conclusions as well as provides avenues of future work.
@@ II.1 Electron-phonon coupling
$g_{nm\nu}=\sqrt{\frac{\hbar}{2M_{0}\omega_{\nu}}}\braket{\psi_{n}|\,\partial U^{\{\tau\}}_{\rm KS}/\partial x_{\nu}\,|\psi_{m}}$ (1)
@@
Figure 2: (a) Schematic representation of the transitions of the PES from the monomorphous to the polymorphous, and from the polymorphous to the reference structure, indicated by ball and stick models of cubic FAPbI3. (b) DFT total energy lowering per formula unit ($\Delta U$) of polymorphous cubic halide perovskites relative to their reference structures, evaluated using Eq. (11), assuming $T=300$ K, and 10 configurations in $2\times 2\times 2$ supercells for each system. One formula unit (f. u.) refers to the unit cell containing 5 atoms. Horizontal gray, red, and light-blue lines indicate the mean $\Delta U$ of FA-, MA-, and Cs-based compounds, respectively. (c-h) Scatter plots showing the dependence of the band gap and effective mass increase due to polymorphism on the average B-X bond length and B-X-B bond angle variations. The 120 data points correspond to 10 polymorphous configurations generated for 4 FA (black), 4 MA (red), and 4 Cs (blue) -based compounds. The straight thick lines represent the least squares regression fit to the data, with the slopes indicated on each plot. The shaded regions represent three standard deviations on either side of the lines. All calculations refer to the DFT-PBEsol approximation. Our calculations for the band gaps and effective masses include spin-orbit coupling (SOC) effects.
@@ II.2 Temperature-dependent electron energies
where the first term inside the square brackets represents the Fan-Migdal contribution and $h_{n\nu\nu}=\hbar/2M_{0}\omega_{\nu}\braket{\psi_{n}|\,\partial^{2}U^{\{\tau\}}_{\rm KS}/{\partial x_{\nu}^{2}}\,|\psi_{n}}$ represents the Debye-Waller contribution to electron-phonon renormalized band structures Giustino (2017). We note that the electron-phonon matrix elements [Eq. (1)] entering Eq. (4), when computed within the harmonic approximation, may be ill-defined if the system exhibits imaginary phonon modes. This is usually an artifact of applying the harmonic approximation to an inherently anharmonic system. This underscores the need to explicitly consider the system’s anharmonicity, as discussed in Secs. II.3 and II.4. Another source of anharmonicity affecting the electronic structure at finite temperatures is thermal lattice expansion, which is treated separately in this work, as discussed in Ref. [Zacharias et al., 2026] and Sec. VI.6.
@@ II.3 Polymorphous structure exploration
Using the example of CsPbI3 for a $2\times 2\times 2$ supercell, we tested the values $x_{\nu}=\{0.05,0.5,1,2,3,4,5\}$ Å  to initially displace the atoms of the reference structure using Eq. (5), followed by structural relaxation. For $x_{\nu}=\{0.5,1,2\}$ Å, the system reached a polymorphous structure, resulting in a total energy lowering of 121 meV relative to the reference structure. For $x_{\nu}=\{3,4,5\}$ Å, the system converged to a saddle point with a total energy reduction of 108 meV, whereas for $x_{\nu}=0.05$ Å, the system returned back to the reference structure.
@@
Here, the summation is performed over stable phonons modes with $\omega^{2}_{\nu}>0$ and $S_{\nu}$ represents signs associated with each eigenvector. These signs are assigned combinatorially by ZG.x of EPW Lee et al. (2023) to ensure that the anisotropic displacement parameters are closely reproduced Zacharias and Giustino (2020). Using CsPbI3, we tested ZG displacements for $T=0$ K. While the system after the escape step moves toward a less stable configuration than the reference structure, the final relaxation step leads to a polymorphous structure with the same energy lowering of 121 meV, as achieved using Eq. (5) for $x_{\nu}=\{0.5,1,2\}$ Å. We note that for generating polymorphous structures with distinct atomic coordinates, one can employ different initial sets of ZG displacements by using various permutations of the set of signs $S_{\nu}$ or different temperatures $T$.